OpenStudy (anonymous):

Express the radical in simplest form: the square root of 40 A. 2 and the square root of 20 B. 4 and the square root of 10 C. 2 and the square root of 10

OpenStudy (wasiqss):

b

OpenStudy (wasiqss):

sorry c

jimthompson5910 (jim_thompson5910):

\[\Large \sqrt{40}\] \[\Large \sqrt{4*10}\] \[\Large \sqrt{4}*\sqrt{10}\] \[\Large 2\sqrt{10}\] So choice C

OpenStudy (anonymous):

right ans is C